# Instability and breaking of internal waves in a horizontal shear layer

### v1

@@ 3.2 Inclined computational domain
The domain length in the periodic $\hat{x}$ and $\hat{z}$ directions is taken to be $L_{x}=L_{z}=8\pi$, whilst a larger extent $y\in(-100,100)$ is used in the direction of shear to accommodate the initially separated wave packet and shear layer. The grid is evenly spaced in the $\hat{x}$ and $\hat{z}$ directions with $N_{x}=N_{z}=648$ grid points in physical space. Polynomial grid stretching is used in the finite difference $y$ direction so that the finest resolution is achieved at the location $y=0$ of maximum shear in the centre of the domain and is equivalent to the grid spacing in the $\hat{x}$ and $\hat{z}$ directions. A total of $N_{y}=1177$ grid points are used in the $y$ direction, with $577$ of these points concentrated in the central region $12.5\leq y\leq 12.5$. The grid resolution is selected such that grid scale remains at most 2.5 times larger than the Kolmogorov length scale $L_{K}=\max(\overline{\varepsilon}Re_{w}^{3})^{-1/4}$ throughout each simulation, where $\overline{\varepsilon}$ is the kinetic energy dissipation rate averaged over $\hat{x}$ and $\hat{z}$ (cf. kaminski2019stratified). Free-slip/no-flux boundary conditions are applied at the walls $y=\pm 100$ ($v=0$, $\partial\phi/\partial y=0$ for $\phi=u,w,b,p$). A sponge layer is employed in regions of width $5$ dimensionless units of length adjacent to the boundaries, within which perturbations are damped to zero quadratically in order to prevent wave reflection back into the centre of the domain.
@@ 3.3 Parameter space
is chosen as $Re_{I}=1200$ for all simulations. Most simulations use an initial wave steepness of $s_{0}=0.75$, though, similar to winters1994three and howland2021shear, we also perform some simulations at a smaller amplitude of $s_{0}=0.25$ to investigate the effect of wave amplitude on the resulting instability. Additional simulations with $\Delta u=0$ (not shown) were performed to verify that the propagating initial wave packet remains stable in the absence of the shear flow. Thus the dynamics of wave instability studied below are instigated by wave-shear interaction and not due solely to parametric self-interaction instability (lombard1996instability) or the interaction of the wave with its amplitude envelope (sutherland2006weakly).
A list of DNS is displayed in table 1. Included are the maximum values of $F_{\mathrm{max}}$ and $s_{\mathrm{max}}$ of the parameter $F$ defined in equation (20) and local steepness $s$ computed along rays using the ray tracing equations outlined above. Larger values of $F_{\mathrm{max}}$ correspond to an increase in the importance of wave advection of streamwise momentum resulting in excess perturbation kinetic energy, whilst values of $s_{\mathrm{max}}>1$ indicate the potential for convective instability. We stress that the precise magnitudes of these parameters are valid only for linear wave perturbations that satisfy the assumptions required for validity of the WKBJ approximation. However, it will be shown that $F_{\mathrm{max}}$ remains useful as a predictor for the partition between kinetic and potential energy in the system even for the strongly nonlinear waves considered here. For the purposes of comparison, we note that in our non-dimensionalisation, simulations 1-4 of staquet2002transport have $\Delta u=4$, $h=4$, $k_{x}=-0.25$, $k_{y}=0.25$, $a=4$, $y_{0}=-8$, $s_{0}=0.3$ and $Re_{w}\approx 2500$.
@@ 4.1 Evolution for $k_{x}<0$ (wave trapping)
Figure 4: $(a)$, $(d)$ and $(g)$ show vertical slices of the perturbation streamwise velocity field for simulations F0.1s0.76(T), F0.1s2.28(T) and F1.0s5.0(T) taken at times corresponding to the panels in the top row of figure 3. Solid black lines are evenly spaced isopycnals (surfaces of constant buoyancy $b=z+\theta$). Mean vertical profiles of the local buoyancy frequency $N^{2}=1+\partial\theta/\partial z$ and vertical shear $S$ averaged in the plane indicated by the dashed magenta lines are shown in the panels to the right. Hatched regions show vertical locations where $N^{2}<0$.
@@
When $s_{0}=0.25$ (first column), distortion of the refracted wave packet is visible in panel $(a)$ at time $t=108$ (approximately 17 dimensionless buoyancy periods). Shortly afterwards, in panel $(b)$, small-scale vorticity structures emerge and propagate rapidly across the trapping level. Disordered small scale motions develop in the center of the shear layer but are quickly damped by viscosity by time $t=150$ in panel $(d)$ at the chosen Reynolds number owing to their relatively small amplitude. Nonetheless, small-scale internal waves, first seen in panel $(c)$, are generated at the edge of the central turbulent region and radiate outward in both directions in a manner qualitatively similar to the results of maffioli2014evolution. For a larger initial wave amplitude $s_{0}=0.75$ but keeping all other parameters the same (second column), instability develops much earlier, and at a considerable distance from the predicted trapping level. Without appealing to the linear theory, there is no obvious indication as to the location of the trapping level from the vorticity snapshots, with a small amount of wave energy clearly able to penetrate through $y=0$. Despite these differences, the subsequent evolution of the instability itself is somewhat similar between the two cases, with the formation of localised vorticity structures that propagate towards and across the center of the shear layer and the generation of internal waves by turbulence. Likewise, the picture of instability remains similar when the shear layer width is decreased by a factor of four and the magnitude of the velocity is increased by a factor of two (third column), though turbulence is clearly most energetic in this case. The qualitative behaviour in this case is also consistent with results obtained by staquet2002transport for comparable initial values of $\Delta u$, $h$ and $k_{x0}$.
@@
Snapshots showing the local wave state prior to breaking are shown in figure 4. Colours show the perturbation streamwise velocity field $u^{\prime}=u-\overline{u{}}$, where an overline represents an average in $x$ and $z$. Isopycnals are included to highlight overturning regions. Local vertical profiles of the buoyancy gradient $N^{2}(z)$ and vertical shear (of the streamwise velocity) $S(z)=|\partial u/\partial z|$ taken at the $y$ location of the maximum velocity amplitude are plotted in the right-hand panels. The results are qualitatively consistent with the predictions from figure 2. In particular, the refraction of the wave by the shear leads to an increase in steepness $s$, as seen from the isopycnals. When the initial amplitude is sufficiently large, local regions of overturning develop (hatched regions in the profiles). It is worth noting the linear theory predicts values of $s>1$ only for parameter values corresponding to the bottom two rows, which is seen to be consistent with the DNS. Larger values of $s$ are naturally associated with increased vertical shear. More importantly for the shear however is the advection of momentum when $\partial U/\partial y$ is large, indicated by the large values of $u^{\prime}$ and corresponding enhanced vertical shear seen in the bottom row of figure 4. This enhanced vertical shear may be associated with the larger value of $F_{\mathrm{max}}$ in simulation F1.0s5.0(T). Notably however, $Ri_{g}<1/4$ only when $N^{2}<0$: there are no regions of the flow susceptible to pure shear instability. This is also consistent with figure 2(b), which shows that although $F$ increases, it never becomes large enough to support a corresponding minimum $Ri_{g}<1/4$ whilst $s<1$.
@@
The three-dimensional behaviour of the wave breaking instability is shown in figure 5 for simulation F0.1s2.28(T). The development of convectively unstable overturns is clear in panel $a)$. At low frequencies, convectively unstable waves break via the production of vorticity that emerges in the form of roll-like structures in the direction parallel to wave crests (howland2021shear). This behaviour has been explained by analysing perturbations to an assumed quasi-steady background state representing the wave at the point of overturning (winters1992instability; parker2021optimal). In general, perturbation energy is extracted from the wave by a mixture of vertical shear production and convection via the buoyancy flux. In this study however, it is less clear that the locally large frequency $\Omega\sim N$ expected in the case of wave trapping satisfies the quasi-steady assumption. Indeed, the vortex structures seen in panels $(e)$ and $(i)$ of figure 3 and in panel $(b)$ of figure 5 appear unique to the horizontal shear scenario. Nonetheless, though the primary instability is fundamentally different to the classical quasi-steady convective wave instability, it is expected that the secondary instability of the streamwise vortices exhibits mixed shear/convection properties (caulfield2000anatomy). The resulting three-dimensionalisation results in the rapid collapse of the convectively unstable region and development of small-scale turbulent structures seen in panel $(c)$.
@@
It is tempting to conclude that turbulent breakdown in the case of wave trapping is ultimately due to a primarily convective instability. However, the case of a smaller-amplitude wave is not consistent with this picture. Figure 6 shows a 3D visualisation of the breaking process for simulation F0.1s0.76(T) with a smaller initial wave amplitude, where it is seen in panel $(a)$ that three-dimensional structure develops prior to the overturning of isopycnals. It will be shown in §5 that wave trapping results in the formation of a narrow velocity jet in the streamwise $x$-direction that is colocated with the region of large isopycnal displacements. Perturbations may then extract energy from the horizontal shear associated with the jet, potentially leading to the generation of vertical vorticity seen in panel $(a)$ and the subsequent breakdown in panel $(b)$. It is worth noting that this behaviour is similar in character to the instability mechanism proposed by wang2020nonlinearii for forced critical layers in a stratified uniform horizontal shear flow.
@@ 4.2 Evolution for $k_{x}\geq 0$
The time evolution of $\zeta_{x}$ for three representative simulations in the case $k_{x}\geq 0$ is shown in figure 7. Note that the left and right columns have parameter values corresponding to the linear calculations shown in figure 2(c)(c) and $(d)$. As expected, the dynamics are fundamentally distinct from the wave trapping scenario explored in the previous section. In the special case $k_{x}=0$ shown in the left-hand column of figure 7, the shear flow does not affect the propagation of the wave so that the slope of phase lines is unchanged in $y$. However, as will shortly be explained, wave advection of mean momentum across the center of the shear layer leads to the local generation of vertical shear. The resulting shear instability creates velocity perturbations which correspond to the small-scale vorticity structures that emerge in the center of the shear layer in panel $(c)$. These local shear instabilities lead to thin layers of turbulence that persist even after the initial wave has passed through the shear layer completely.
@@
Figure 8: The same as figure 4, for simulations (a)(a)-$(c)$ F6.0s0.75_kx0; (d)(d)-$(f)$ F6.0s0.25; (g)(g)-$(i)$ F6.0s0.75. In the right-hand panels, dotted hatching indicates shear-unstable regions where $0\leq Ri_{g}<1/4$, and line hatching shows vertical locations where $N^{2}<0$, as before.
The influence of the advection of background shear by the wave is made abundantly clear by plotting the streamwise perturbation velocity $u^{\prime}$ in figure 8. Horizontal layers of positive and negative $u^{\prime}$ are seen in all three cases, with shear magnitudes $O(1)$ and above. In particular, for the large-amplitude initial conditions, $u^{\prime}$ may be sufficiently large such that regions with $0\leq Ri_{g}<0.25$ develop (top and bottom rows), paving the way for the shear instabilities seen in figure 7(c,k)(c,k). This behaviour is in accordance with the ray-tracing predictions of figure 2(c,d)(c,d). In particular, the linear theory correctly predicts values of $F$ large enough such that $Ri_{g}<0.25$ in simulations F6.0s0.75 and F6.0s0.75_kx0. However, a notable difference from the theory arises from the interference of the incident and reflected waves in simulation F6.0s0.75, leading to narrow regions of local static instability with $N^{2}<0$. As will be seen, this feature plays a prominent role in the triggering of secondary instabilities, though the subsequent development of turbulence and mixing remains very much driven by the enhanced vertical shear.
@@
A three-dimensional visualisation of simulation F6.0s0.75 is shown in figure 9. The interaction of the reflected and incident waves at the center of the shear layer leads to the isopycnal overturning seen in panel $(a)$. Panel $(b)$ shows this overturning region spreading out across the shear layer and in particular highlights the wave advection of background flow vertical vorticity that leads to the vertical shear generation. At the same time, streak-like $\zeta_{z}$ perturbations have emerged in the center of the overturning region. These instabilities evolve to have billow-like structures seen in panel $(c)$ closely resembling those of classical shear instability, where vortices are oriented in the horizontal direction at an oblique angle to the $x$-direction of the background flow. Note that this orientation is not consistent with the prediction $\beta=\pi/2$ from (22), that is, that the most unstable direction is the streamwise direction (and thus the billow axis should be in the $y$-direction). This is likely due to the continuous shearing of the growing perturbations by the background flow. In fact, a similar phenomenon was also reported by lelong1998inertiaa for shear instabilities in convectively stable inertia-gravity waves, where the discrepancy between theory and simulations was attributed to the lack of scale separation between the timescales of wave propagation and the growth rate of instability assumed when appealing to a local gradient Richardson number criterion for instability. The development of small-scale turbulent structures seen in panel $(c)$ is due to a combination of convective instability of the overturning region seen in $(b)$ and secondary instabilities of the developing Kelvin-Helmholtz billows. As such, the precise realisation of the turbulent transition is likely to be sensitive to $Re_{w}$ (and indeed $Pr$).
Further evidence for behaviour dominated by vertical shear instability is provided by simulation F6.0s0.75_kx0, which does not exhibit the wave reflection and subsequent isopycnal overturns seen in simulation F6.0s0.75. A three-dimensional picture of developing secondary instabilities for F6.0s0.75_kx0 is shown in figure 10. The dynamics are similar to those in figure 9, with perturbations oriented at an oblique angle to the $x$-direction. In this case, the angle $\beta$ is even further from the predicted value of $\pi/2$, possibly due to the slower growth rate of instability. As in figure 9, turbulence in figure 10 follows from a subsequent instability of these perturbations. In this case the secondary instabilities seen in panel $(b)$ resemble the instability due to misaligned vortex cores studied by fritts2022multi.

### v2

@@ 3.2 Inclined computational domain
The domain length in the periodic $\hat{x}$ and $\hat{z}$ directions is taken to be $L_{x}=L_{z}=8\pi$, whilst a larger extent $y\in(-100,100)$ is used in the direction of shear to accommodate the initially separated wave packet and shear layer. The grid is evenly spaced in the $\hat{x}$ and $\hat{z}$ directions with $N_{x}=N_{z}=648$ grid points in physical space. Polynomial grid stretching is used in the finite difference $y$ direction so that the finest resolution is achieved at the location $y=0$ of maximum shear in the centre of the domain and is equivalent to the grid spacing in the $\hat{x}$ and $\hat{z}$ directions. A total of $N_{y}=1177$ grid points are used in the $y$ direction, with $577$ of these points concentrated in the central region $12.5\leq y\leq 12.5$. The grid resolution is selected such that grid scale remains at most 2.5 times larger than the Kolmogorov length scale $L_{K}=\max(\overline{\varepsilon}Re_{w}^{3})^{-1/4}$ throughout each simulation, where $\overline{\varepsilon}$ is the kinetic energy dissipation rate averaged over $\hat{x}$ and $\hat{z}$ (cf. Kaminski and Smyth, 2019). Free-slip/no-flux boundary conditions are applied at the walls $y=\pm 100$ ($v=0$, $\partial\phi/\partial y=0$ for $\phi=u,w,b,p$). A sponge layer is employed in regions of width $5$ dimensionless units of length adjacent to the boundaries, within which perturbations are damped to zero quadratically in order to prevent wave reflection back into the centre of the domain.
@@ 3.3 Parameter space
is chosen as $Re_{I}=1200$ for all simulations. Most simulations use an initial wave steepness of $s_{0}=0.75$, though, similar to Winters and D’Asaro (1994) and Howland et al. (2021), we also perform some simulations at a smaller amplitude of $s_{0}=0.25$ to investigate the effect of wave amplitude on the resulting instability. Additional simulations with $\Delta u=0$ (not shown) were performed to verify that the propagating initial wave packet remains stable in the absence of the shear flow. Thus the dynamics of wave instability studied below are instigated by wave-shear interaction and not due solely to parametric self-interaction instability (Lombard and Riley, 1996a) or the interaction of the wave with its amplitude envelope (Sutherland, 2006b).
A list of DNS is displayed in table 1. Included are the maximum values of $F_{\mathrm{max}}$ and $s_{\mathrm{max}}$ of the parameter $F$ defined in equation (20) and local steepness $s$ computed along rays using the ray tracing equations outlined above. Larger values of $F_{\mathrm{max}}$ correspond to an increase in the importance of wave advection of streamwise momentum resulting in excess perturbation kinetic energy, whilst values of $s_{\mathrm{max}}>1$ indicate the potential for convective instability. We stress that the precise magnitudes of these parameters are valid only for linear wave perturbations that satisfy the assumptions required for validity of the WKBJ approximation. However, it will be shown that $F_{\mathrm{max}}$ remains useful as a predictor for the partition between kinetic and potential energy in the system even for the strongly nonlinear waves considered here. For the purposes of comparison, we note that in our non-dimensionalisation, simulations 1-4 of Staquet and Huerre (2002) have $\Delta u=4$, $h=4$, $k_{x}=-0.25$, $k_{y}=0.25$, $a=4$, $y_{0}=-8$, $s_{0}=0.3$ and $Re_{w}\approx 2500$.
@@ 4.1 Evolution for $k_{x}<0$ (wave trapping)
Figure 4: $(a)$, $(d)$ and $(g)$ show vertical slices of the perturbation streamwise velocity field for simulations F0.1s0.76(T), F0.1s2.28(T) and F1.0s5.0(T) taken at times corresponding to the panels in the top row of figure 3. Solid black lines are evenly spaced isopycnals (surfaces of constant buoyancy $b=z+\theta$). Vertical profiles of the local buoyancy frequency $N^{2}=1+\partial\theta/\partial z$ and vertical shear $S$ are shown in the panels to the right. The locations at which these profiles are taken are indicated by the dashed magenta lines. Hatched regions show vertical locations where $N^{2}<0$.
@@
When $s_{0}=0.25$ (first column), distortion of the refracted wave packet is visible in panel $(a)$ at time $t=108$ (approximately 17 dimensionless buoyancy periods). Shortly afterwards, in panel $(b)$, small-scale vorticity structures emerge and propagate rapidly across the trapping level. Disordered small scale motions develop in the center of the shear layer but are quickly damped by viscosity by time $t=150$ in panel $(d)$ at the chosen Reynolds number owing to their relatively small amplitude. Nonetheless, small-scale internal waves, first seen in panel $(c)$, are generated at the edge of the central turbulent region and radiate outward in both directions in a manner qualitatively similar to the results of Maffioli et al. (2014). For a larger initial wave amplitude $s_{0}=0.75$ but keeping all other parameters the same (second column), instability develops much earlier, and at a considerable distance from the predicted trapping level. Without appealing to the linear theory, there is no obvious indication as to the location of the trapping level from the vorticity snapshots, with a small amount of wave energy clearly able to penetrate through $y=0$. Despite these differences, the subsequent evolution of the instability itself is somewhat similar between the two cases, with the formation of localised vorticity structures that propagate towards and across the center of the shear layer and the generation of internal waves by turbulence. Likewise, the picture of instability remains similar when the shear layer width is decreased by a factor of four and the magnitude of the velocity is increased by a factor of two (third column), though turbulence is clearly most energetic in this case. The qualitative behaviour in this case is also consistent with results obtained by Staquet and Huerre (2002) for comparable initial values of $\Delta u$, $h$ and $k_{x0}$.
@@
Snapshots showing the local wave state prior to breaking are shown in figure 4. Colours show the perturbation streamwise velocity field $u^{\prime}=u-\overline{u{}}$, where an overline represents an average in $x$ and $z$. Isopycnals are included to highlight overturning regions. Local vertical profiles of the buoyancy gradient $N^{2}(z)$ and vertical shear (of the streamwise velocity) $S(z)=|\partial u/\partial z|$ taken at the $y$ location of the maximum velocity amplitude are plotted in the right-hand panels. The results are qualitatively consistent with the predictions from figure 2. In particular, the refraction of the wave by the shear leads to an increase in steepness $s$, as seen from the isopycnals. When the initial amplitude is sufficiently large, local regions of overturning develop (hatched regions in the profiles). It is worth noting the linear theory predicts values of $s>1$ only for parameter values corresponding to the bottom two rows, which is seen to be consistent with the DNS. Larger values of $s$ are naturally associated with increased vertical shear. More important for the shear however is the advection of momentum when $\partial U/\partial y$ is large, indicated by the large values of $u^{\prime}$ and corresponding enhanced vertical shear seen in the bottom row of figure 4. This enhanced vertical shear may be associated with the larger value of $F_{\mathrm{max}}$ in simulation F1.0s5.0(T). Notably however, $Ri_{g}<1/4$ only when $N^{2}<0$: there are no regions of the flow susceptible to pure shear instability. This is also consistent with figure 2(b), which shows that although $F$ increases, it never becomes large enough to support a corresponding minimum $Ri_{g}<1/4$ whilst $s<1$.
@@
The three-dimensional behaviour of the wave breaking instability is shown in figure 5 for simulation F0.1s2.28(T). The development of convectively unstable overturns is clear in panel $a)$. At low frequencies, convectively unstable waves break via the production of vorticity that emerges in the form of roll-like structures in the direction parallel to wave crests (Howland et al., 2021). This behaviour has been explained by analysing perturbations to an assumed quasi-steady background state representing the wave at the point of overturning (Winters and Riley, 1992; Parker et al., 2021). In general, perturbation energy is extracted from the wave by a mixture of vertical shear production and convection via the buoyancy flux. In this study however, it is less clear that the locally large frequency $\Omega\sim N$ expected in the case of wave trapping satisfies the quasi-steady assumption. Indeed, the vortex structures seen in panels $(e)$ and $(i)$ of figure 3 and in panel $(b)$ of figure 5 appear unique to the horizontal shear scenario. Nonetheless, though the primary instability is fundamentally different to the classical quasi-steady convective wave instability, it is expected that the secondary instability of the streamwise vortices exhibits mixed shear/convection properties (Caulfield and Peltier, 2000). The resulting three-dimensionalisation results in the rapid collapse of the convectively unstable region and development of small-scale turbulent structures seen in panel $(c)$.
@@
It is tempting to conclude that turbulent breakdown in the case of wave trapping is ultimately due to a primarily convective instability. However, the case of a smaller-amplitude wave is not consistent with this picture. Figure 6 shows a 3D visualisation of the breaking process for simulation F0.1s0.76(T) with a smaller initial wave amplitude, where it is seen in panel $(a)$ that three-dimensional structure develops prior to the overturning of isopycnals. It will be shown in §5 that wave trapping results in the formation of a narrow velocity jet in the streamwise $x$-direction that is colocated with the region of large isopycnal displacements. Perturbations may then extract energy from the horizontal shear associated with the jet, potentially leading to the generation of vertical vorticity seen in panel $(a)$ and the subsequent breakdown in panel $(b)$. It is worth noting that this behaviour is similar in character to the instability mechanism proposed by Wang and Balmforth (2021) for forced critical layers in a stratified uniform horizontal shear flow.
@@ 4.2 Evolution for $k_{x}\geq 0$
The time evolution of $\zeta_{x}$ for three representative simulations in the case $k_{x}\geq 0$ is shown in figure 7 (simulations F3.0s0.25 and F3.0s0.75 are qualitatively similar to F6.0s0.25 and F6.0s0.75). Note that the left and right columns have parameter values corresponding to the linear calculations shown in figure 2(c)(c) and $(d)$. As expected, the dynamics are fundamentally distinct from the wave trapping scenario explored in the previous section. In the special case $k_{x}=0$ shown in the left-hand column of figure 7, the shear flow does not affect the propagation of the wave so that the slope of phase lines is unchanged in $y$. However, as will shortly be explained, wave advection of mean momentum across the center of the shear layer leads to the local generation of vertical shear. The resulting shear instability creates velocity perturbations which correspond to the small-scale vorticity structures that emerge in the center of the shear layer in panel $(c)$. These local shear instabilities lead to thin layers of turbulence that persist even after the initial wave has passed through the shear layer completely. Curiously, there are noticeable variations in the development of shear instabilities at different vertical heights (most clearly seen around $z\approx 20$). Vorticity perturbations have a sufficiently large vertical extent such that it seems reasonable to suppose that the instabilities developing in one layer influence the dynamics of adjacent layers. We speculate that interactions between layers could quickly magnify initially small differences in the timing of instability onset in neighbouring layers that occur due to random fluctuations in the background noise field.
@@
Figure 8: The same as figure 4, for simulations (a)(a)-$(c)$ F6.0s0.75_kx0; (d)(d)-$(f)$ F6.0s0.25; (g)(g)-$(i)$ F6.0s0.75, where snapshots are taken at times corresponding to the top row panels in figure 7. In the right-hand panels, dotted hatching indicates shear-unstable regions where $0\leq Ri_{g}<1/4$, and line hatching shows vertical locations where $N^{2}<0$, as before.
The influence of the advection of background shear by the wave is made abundantly clear by plotting the streamwise perturbation velocity $u^{\prime}$ in figure 8. Horizontal layers of positive and negative $u^{\prime}$ are seen in all three cases, with shear magnitudes $O(1)$ and above. In particular, for the large-amplitude initial conditions, $u^{\prime}$ may be sufficiently large such that regions with $0<Ri_{g}<0.25$ develop (top and bottom rows), whose locations correspond to the layers of shear instabilities seen developing in figure 7(c,k)(c,k). This behaviour is in accordance with the ray-tracing predictions of figure 2(c,d)(c,d). Specifically, the linear theory correctly predicts values of $F$ large enough such that $Ri_{g}<0.25$ in simulations F6.0s0.75 and F6.0s0.75_kx0, whereas for simulation F6.0s0.25 (which does not become turbulent), $\min(Ri_{g})\approx(1-s_{\mathrm{max}})/F_{\mathrm{max}}^{2}s_{\mathrm{max}}^{2}\approx 0.42>0.25$. On the other hand, a notable difference from the theory arises from the interference of the incident and reflected waves in simulation F6.0s0.75, leading to narrow regions of local static instability with $N^{2}<0$ in locations where the instabilities seen in figure 7(j)(j) develop. Clearly in this case, convective instabilities develop on a shorter time scale than shear instabilities in the adjacent regions with $0<Ri_{g}<0.25$. However, though regions with $N^{2}<0$ play a dominant role in the triggering of secondary instabilities, it will be shown that the subsequent development of turbulence and mixing remains very much driven by the enhanced vertical shear.
@@
A three-dimensional visualisation of simulation F6.0s0.75 is shown in figure 9. The interaction of the reflected and incident waves at the center of the shear layer leads to the isopycnal overturning seen in panel $(a)$. Panel $(b)$ shows this overturning region spreading out across the shear layer and in particular highlights the wave advection of background flow vertical vorticity that leads to the vertical shear generation. At the same time, streak-like $\zeta_{z}$ perturbations have emerged in the center of the overturning region. These instabilities evolve to have billow-like structures seen in panel $(c)$ closely resembling those of classical shear instability, where vortices are oriented in the horizontal direction at an oblique angle to the $x$-direction of the background flow. Note that this orientation is not consistent with the prediction $\beta=\pi/2$ from (22), that is, that the most unstable direction is the streamwise direction (and thus the billow axis should be in the $y$-direction). This is likely due to the continuous shearing of the growing perturbations by the background flow. In fact, a similar phenomenon was also reported by Lelong and Dunkerton (1998a) for shear instabilities in convectively stable inertia-gravity waves, where the discrepancy between theory and simulations was attributed to the lack of scale separation between the timescales of wave propagation and the growth rate of instability assumed when appealing to a local gradient Richardson number criterion for instability. The development of small-scale turbulent structures seen in panel $(c)$ is due to a combination of convective instability of the overturning region seen in $(b)$ and secondary instabilities of the developing Kelvin-Helmholtz billows. As such, the precise realisation of the turbulent transition is likely to be sensitive to $Re_{w}$ (and indeed $Pr$).
Further evidence for behaviour dominated by vertical shear instability is provided by simulation F6.0s0.75_kx0, which does not exhibit the wave reflection and subsequent isopycnal overturns seen in simulation F6.0s0.75. A three-dimensional picture of developing secondary instabilities for F6.0s0.75_kx0 is shown in figure 10. The dynamics are similar to those in figure 9, with perturbations oriented at an oblique angle to the $x$-direction. In this case, the angle $\beta$ is even further from the predicted value of $\pi/2$, possibly due to the slower growth rate of instability. As in figure 9, turbulence in figure 10 follows from a subsequent instability of these perturbations. In this case the secondary instabilities seen in panel $(b)$ resemble the instability due to misaligned vortex cores studied by Fritts et al. (2022).